OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. r times the square root of r − 3 times r times the square root of r + 4 times r times the square root of r 2r 2 times r times the square root of r 2 the square root of r the square root of r

OpenStudy (agent0smith):

You'd be better off learning how to use the equation editor... that way people don't have to guess at what your wordy expressions are \[\Large r \sqrt {r-3} *r \sqrt{r+4} * r * \sqrt r\] I have no idea, to be honest... "r times the square root of r − 3 times r times the square root of r + 4 times r times the square root of r" this is way too hard to interpret.

OpenStudy (anonymous):

\[r \sqrt{r}-3r \sqrt{r}+4r \sqrt{r}\]

OpenStudy (anonymous):

combine like terms or factor out the \(\sqrt{r}\) same thing really

OpenStudy (agent0smith):

Thank you. Make sure you use it everytime from now on. Just combine like terms or factor as satellite said... they all have a sqrt r. \[\Large r \sqrt{r}-3r \sqrt{r}+4r \sqrt{r} = \sqrt r (r - 3r +4r)\]

OpenStudy (anonymous):

\[2r \sqrt{r}?\]

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

thank you! can you help me with one more?

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. \[\sqrt{16y^3}-\sqrt{y^3}\]

OpenStudy (anonymous):

a) \[3y \sqrt{y} b) 3y^2\sqrt{y} c) 4y^2\sqrt{y} d) 4y \sqrt{y}\]

OpenStudy (agent0smith):

\[\Large \sqrt{16y^3}-\sqrt{y^3} = \sqrt{16*y^2 *y}-\sqrt{y^2*y}\] \[\Large \sqrt{16} \sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y}\]

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

the square root of 16 is 4

OpenStudy (agent0smith):

Simplify everything here as much as you can \[\Large \sqrt{16} \sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y}\] you know sqrt16 is 4, so \[\Large 4\sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y}\]What else can you simplify?

OpenStudy (anonymous):

\[\sqrt{y^2}?\]

OpenStudy (agent0smith):

Yes, what's the sqrt of y^2?

OpenStudy (anonymous):

\[4\sqrt{y^2}-\sqrt{y^2}= 3\sqrt{y^2}\] am I right?

OpenStudy (agent0smith):

Not quite... go back to this: \[\Large 4\sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y} \] What's the sqrt of y^2?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

y?

OpenStudy (agent0smith):

Think of it this way... Whats the sqrt of 4^2?

OpenStudy (anonymous):

2

OpenStudy (agent0smith):

No... 4^2 = 16. What's the sqrt of 16?

OpenStudy (anonymous):

4

OpenStudy (agent0smith):

So the sqrt of 4^2 is 4. What's the sqrt of y^2, then?

OpenStudy (anonymous):

2y? ugh im confused :(

OpenStudy (agent0smith):

Well notice the sqrt of 4^2 is 4. What's the sqrt of 5^2? what's the sqrt of 2^2?

OpenStudy (anonymous):

5,2

OpenStudy (agent0smith):

Notice the pattern? Sqrt of 4^2 = 4. Sqrt 5^2 = 5. Sqrt 2^2 = 2. Sqrt y^2 = ...?

OpenStudy (anonymous):

y

OpenStudy (agent0smith):

Excellent! What about the sqrt of x^2?

OpenStudy (anonymous):

x

OpenStudy (agent0smith):

Good :) Now simplify \[\Large 4\sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y}\]

OpenStudy (anonymous):

\[4y \sqrt{y}\]

OpenStudy (agent0smith):

Not quite but you're on the right track\[\Large 4\sqrt {y^2} \sqrt {y }-\sqrt{y^2} \sqrt {y}=\] replace the two sqrt y^2's with y \[\Large 4y \sqrt {y }-y \sqrt {y}\]

OpenStudy (anonymous):

\[3y \sqrt{y}\] so it's a

OpenStudy (agent0smith):

I think so, i didn't pay attention to the answer options. But good job!

OpenStudy (anonymous):

thank you!